OpenStudy (studybird123):

A circle has a radius of 10 yards. What is the circumference of the circle? (Express your answer using π.) A. 5π yards B. 10π yards C. 20π yards D. 30π

OpenStudy (anonymous):

ok, do you know you're formulas for circumference?

OpenStudy (anonymous):

\[C=2\pi r\]

OpenStudy (studybird123):

um is it 3.14* diameter?

OpenStudy (anonymous):

yes pi*d is one of them 2*pi*r is the other

OpenStudy (studybird123):

I keep getting 62.83

OpenStudy (anonymous):

1 sec, let me get a piece of paper

OpenStudy (anonymous):

Because you have to express your answer in pi, just do 2*10=20, and 20*pi=62.83 So the answer is 20 pi

OpenStudy (anonymous):

well yea thats correct :P

OpenStudy (studybird123):

ohh thats how you do it

OpenStudy (anonymous):

nvm its not

OpenStudy (mathstudent55):

They tell you to leave the answer in terms of pi. Don't multiply by 3.14. Just write the answer as the diameter (20) time pi. \(C = 20 \pi\)

OpenStudy (anonymous):

oh yes it is,

OpenStudy (studybird123):

can you guys like give each other medals since i can give only one person thanks!

OpenStudy (anonymous):

pfffft

OpenStudy (anonymous):

I gave one to mathstudent

OpenStudy (studybird123):

ok thanks orangemaste and jacob for the medal

OpenStudy (anonymous):

good luck study bird

OpenStudy (mathstudent55):

i gave one to orangemaster

OpenStudy (studybird123):

thanks mathstudent

OpenStudy (mathstudent55):

You're welcome.